1/2x + 2/5 (x-2) = 1
Answer by checkley77(12844)

1/2x+2/5(x-2)=1
X/2+2(X-2)/5=1
X/2+(2X-4)/5=1
[(5X)+2(2X-4)]/2*5=1
[5X+4X-8]/10=1
(9X-8)/10=1
9X-8=10
9X=10+8
9X=18
X=18/9
X=2 ANSWER.
PROOF:
2/2+2(2-2)/5=1
1+2*0/5=1
1+0=1
1=1